CFFI-SYS Interface Specification

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005-2006, James Bielman <jamesjb at jamesjb.com> Permission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the “Software”), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, merge, publish, distribute, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions: The above copyright notice and this permission notice shall be included in all copies or substantial portions of the Software. The software is provided “as is”, without warranty of any kind, express or implied, including but not limited to the warranties of merchantability, fitness for a particular purpose and noninfringement. In no event shall the authors or copyright holders be liable for any claim, damages or other liability, whether in an action of contract, tort or otherwise, arising from, out of or in connection with the software or the use or other dealings in the software.

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Built-In Foreign Types . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Operations on Built-in Foreign Types . . . . . . . . . 3 4 Basic Pointer Operations . . . . . . . . . . . . . . . . . . . . . . . 4 5 Foreign Memory Allocation . . . . . . . . . . . . . . . . . . . . 5 6 Memory Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7 Foreign Function Calling. . . . . . . . . . . . . . . . . . . . . . . . 7 8 Loading Foreign Libraries. . . . . . . . . . . . . . . . . . . . . . . 8 9 Foreign Globals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Symbol Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

1 Introduction

CFFI, the Common Foreign Function Interface, purports to be a portable foreign function interface for Common Lisp.

This specification defines a set of low-level primitives that must be defined for each Lisp implementation supported by CFFI. These operators are defined in the CFFI-SYS package.

The CFFI package uses the CFFI-SYS interface to implement an extensible foreign type system with support for typedefs, structures, and unions, a declarative interface for defining foreign function calls, and automatic conversion of foreign function arguments to/from Lisp types.

Please note the following conventions that apply to everything in CFFI-SYS:

• Functions in CFFI-SYS that are low-level versions of functions exported from the CFFI package begin with a leading percent-sign (eg. %mem-ref).

• Where “foreign type” is mentioned as the kind of an argument, the meaning is restricted to that subset of all foreign types defined in Chapter 2 [Built-In Foreign Types], page 2. Support for higher-level types is always defined in terms of those lower-level types in CFFI proper.

2 Built-In Foreign Types

[Foreign Type] :char

[Foreign Type] :unsigned-char

[Foreign Type] :short

[Foreign Type] :unsigned-short

[Foreign Type] :int

[Foreign Type] :unsigned-int

[Foreign Type] :long

[Foreign Type] :unsigned-long

[Foreign Type] :long-long

[Foreign Type] :unsigned-long-long

These types correspond to the native C integer types according to the ABI of the system the Lisp implementation is compiled against.

[Foreign Type] :int8 [Foreign Type] :uint8 [Foreign Type] :int16 [Foreign Type] :uint16 [Foreign Type] :int32 [Foreign Type] :uint32 [Foreign Type] :int64 [Foreign Type] :uint64

Foreign integer types of specific sizes, corresponding to the C types defined in

stdint.h.

[Foreign Type] :size [Foreign Type] :ssize [Foreign Type] :ptrdiff [Foreign Type] :time

Foreign integer types corresponding to the standard C types (without the _t suffix).

Implementor’s note: I’m sure there are more of these that could be useful, let’s add any types that can’t be defined portably to this list as necessary.

[Foreign Type] :float [Foreign Type] :double

The :float type represents a C float and a Lisp single-float. :double represents

a C double and a Lisp double-float.

[Foreign Type] :pointer

A foreign pointer to an object of any type, corresponding to void *.

[Foreign Type] :void

No type at all. Only valid as the return type of a function.

3 Operations on Built-in Foreign Types

[Function] %foreign-type-size type ⇒ size

Return the size, in bytes, of objects having foreign type type. An error is signalled if type is not a known built-in foreign type.

[Function] %foreign-type-alignment type ⇒ alignment

Return the default alignment in bytes for structure members of foreign type type. An error is signalled if type is not a known built-in foreign type. Implementor’s note: Maybe this should take an optional keyword argument specifying an alternate alignment system, eg. :mac68k for 68000-compatible alignment on Darwin.

4 Basic Pointer Operations

[Function] pointerp ptr ⇒ boolean

Return true if ptr is a foreign pointer.

[Function] null-pointer ⇒ pointer

Return a null foreign pointer.

[Function] null-pointer-p ptr ⇒ boolean

Return true if ptr is a null foreign pointer.

[Function] make-pointer address ⇒ pointer

Return a pointer corresponding to the numeric integer address.

[Function] inc-pointer ptr offset ⇒ pointer

Return the result of numerically incrementing ptr by offset.

5 Foreign Memory Allocation

[Function] foreign-alloc size ⇒ pointer

Allocate size bytes of foreign-addressable memory and return a pointer to the allocated block. An implementation-specific error is signalled if the memory cannot be allocated.

[Function] foreign-free ptr ⇒ unspecified

Free a pointer ptr allocated by foreign-alloc. The results are undefined if ptr is used after being freed.

[Macro] with-foreign-pointer (var size &optional size-var) &body body

Bind var to a pointer to size bytes of foreign-accessible memory during body. Both ptr and the memory block it points to have dynamic extent and may be stack allocated if supported by the implementation. If size-var is supplied, it will be bound to size during body.

6 Memory Access

[Accessor] %mem-ref ptr type &optional offset

Dereference a pointer offset bytes from ptr to an object for reading (or writing when used with setf) of built-in type type.

Example

;; An impractical example, since time returns the time as well,
;; but it demonstrates %MEM-REF. Better (simple) examples wanted!
(with-foreign-pointer (p (foreign-type-size :time))
  (foreign-funcall "time" :pointer p :time)
  (%mem-ref p :time))

7 Foreign Function Calling

[Macro] %foreign-funcall name {arg-type arg}* &optional result-type ⇒ object

[Macro] %foreign-funcall-pointer ptr {arg-type arg}* &optional result-type

⇒ object

Invoke a foreign function called name in the foreign source code. Each arg-type is a foreign type specifier, followed by arg, Lisp data to be converted to foreign data of type arg-type. result-type is the foreign type of the function’s return value, and is assumed to be :void if not supplied. %foreign-funcall-pointer takes a pointer ptr to the function, as returned by foreign-symbol-pointer, rather than a string name.

[Macro] %foreign-funcall-varargs name ({fixed-type arg}*) {vararg-type arg}* &optional result-type ⇒ object

[Macro] %foreign-funcall-varargs-pointer ptr ({fixed-type arg}*)

{vararg-type arg}* &optional result-type ⇒ object

Invoke a foreign variadic function called name in the foreign source code. Each fixed-type and vararg-type is a foreign type specifier, followed by arg, Lisp data to be converted to foreign data of type arg-type. result-type is the foreign type of the function’s return value, and is assumed to be :void if not supplied. %foreign-funcall-pointer-varargs takes a pointer ptr to the variadic function, as returned by foreign-symbol-pointer, rather than a string name. Both functions have default implementation which call %foreign-funcall and

%foreign-funcall-pointer approprietly.

Examples

;; Calling a standard C library function:
(%foreign-funcall "sqrtf" :float 16.0 :float) ⇒ 4.0
;; Dynamic allocation of a buffer and passing to a function:
(with-foreign-ptr (buf 255 buf-size)
  (%foreign-funcall "gethostname" :pointer buf :size buf-size :int)
  ;; Convert buf to a Lisp string using MAKE-STRING and %MEM-REF or
  ;; a portable CFFI function such as CFFI:FOREIGN-STRING-TO-LISP.
  )

8 Loading Foreign Libraries

[Function] %load-foreign-library name ⇒ unspecified

Load the foreign shared library name. Implementor’s note: There is a lot of behavior to decide here. Currently I lean toward not requiring NAME to be a full path to the library so we can search the system library directories (maybe even get LD LIBRARY PATH from the environment) as necessary.

9 Foreign Globals

[Function] foreign-symbol-pointer name ⇒ pointer

Return a pointer to a foreign symbol name.

Symbol Index

%

%foreign-funcall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

%foreign-funcall-pointer . . . . . . . . . . . . . . . . . . . . . 7
%foreign-funcall-varargs . . . . . . . . . . . . . . . . . . . . . 7
%foreign-funcall-varargs-pointer . . . . . . . . . . . . 7
%foreign-type-alignment . . . . . . . . . . . . . . . . . . . . . . 3

%foreign-type-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 %load-foreign-library . . . . . . . . . . . . . . . . . . . . . . . . 8 %mem-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

:

:char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :float. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :int16. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :int32. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :int64. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :int8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :long-long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :ptrdiff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :short. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :ssize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :uint16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :uint32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :uint64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :uint8. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :unsigned-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :unsigned-int. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :unsigned-long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :unsigned-long-long. . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :unsigned-short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 :void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

F

foreign-alloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 foreign-free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 foreign-symbol-pointer . . . . . . . . . . . . . . . . . . . . . . . 9

I

inc-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

M

make-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

N

null-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 null-pointer-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

P

pointerp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

W

with-foreign-pointer. . . . . . . . . . . . . . . . . . . . . . . . . . 5